OpenStudy (anonymous):

√50+√2 Which statement is true about the expression? a. It is irrational and equal to 2√13 b. It is irrational and equal to 6√2 c. It is rational and equal to 2. d. It is rational and equal to 6.

OpenStudy (anonymous):

\[ \sqrt{50} + \sqrt{2} = \sqrt{2 \cdot 25} + \sqrt{2} = \sqrt{2} \cdot \sqrt{25} + \sqrt{2} = \\ = 5 \cdot \sqrt{2} + \sqrt{2} = 6 \cdot \sqrt{2} \]